OpenStudy (anonymous):

Will give a medal... Please help me... What is the solution of log2x − 5 (25) = 2? x = −5 x = 2 x = 5 x = 3

OpenStudy (owlcoffee):

is it base or is it multiplying?

OpenStudy (anonymous):

2x-5 is the base

OpenStudy (owlcoffee):

okay, logarithm is defined the following way: \[\log_{b} a=c <=> b ^{c}=a\] we can read it like this: "The logarithm with base 'b' of a number 'a' and equal a number 'c' is true only if 'b' to the power of 'c' is equal to 'a'" applying that to your problem: \[\log_{2x-5} 25=2\] applying definition: \[\log_{2x-5} 25 = 2 <=>(2x-5)^{2}=25\] so, traducing it, we wiped out the logarithm and made it into a one-variable equality: \[(2x-5)^{2}=25\] All you have to do is solve for "x"

OpenStudy (anonymous):

So, X= -5? If I am correct.

OpenStudy (owlcoffee):

REmember, when you apply square root it's always plus/minus.

OpenStudy (anonymous):

hey did you get this one right

OpenStudy (owlcoffee):

I think she/he did.